OpenStudy (anonymous):

find d^2y/dx^2 of x^2 + y^2 = 7

OpenStudy (phi):

can you first find dy/dx ?

OpenStudy (anonymous):

yeah x/y

OpenStudy (phi):

dy/dx = - x/y

OpenStudy (anonymous):

yeah i fixed what i did a second ago. i just dont know how to get the second part

OpenStudy (phi):

as a first step, take the derivative of both sides of \[ y' = -\frac{x}{y} \]

OpenStudy (anonymous):

so the derivative of -x and -y?

OpenStudy (phi):

use the quotient rule or rewrite as -1*x*y^(-1) and use the product rule

OpenStudy (anonymous):

i still got -x/y

OpenStudy (phi):

yes, now that we have \[ \frac{dy}{dx} = - \frac{x}{y} \] now take the derivative a second time: \[ \frac{d^2y}{dx^2} = -\frac{d}{dx}\left( \frac{x}{y}\right) \]

OpenStudy (phi):

quotient rule: \[ d \frac{u}{v} = \frac{ v\ du - u\ dv}{v^2} \]

OpenStudy (anonymous):

so the answer is -x^2+y2/y^2?

OpenStudy (phi):

what did you get for d (-x/y) ?

OpenStudy (anonymous):

i dont know... you said what the formula was and that was the answer choice that was closest

OpenStudy (phi):

oh we have to do this problem in steps. even after you find d (-x/y) there is one more step after that to get a "simplified" version of d^2 y/dx^2 you need to find d(-x/y)

OpenStudy (anonymous):

ok to be honest i dont understand implicit differentistion at all

OpenStudy (phi):

take 5 minutes and look over this post from yesterday http://openstudy.com/users/thatguyintheback#/updates/52b4e4aae4b02932bbbfdb0b then we will pick up here.

OpenStudy (phi):

I assume if you had the problem \[ \frac{d}{dx} \left(\frac{x}{\sin x}\right) \] you could use the quotient rule ? it is the same for \[ \frac{d}{dx} \left(\frac{x}{y}\right) \] we will use the fact that \( \frac{d}{dx} y= \frac{dy}{dx} \)

OpenStudy (anonymous):

ok i have to eat breakfast but ill be on later and try to understand. ill work on it when im gone. thanks for your help!

OpenStudy (phi):

ok, meanwhile \[- \frac{d}{dx}\left(\frac{x}{y}\right) = -\left(\frac{ y \frac{d}{dx}x - x \frac{d}{dx}y}{y^2} \right) \\ -\left(\frac{ y - x \frac{dy}{dx}}{y^2} \right) \\ \left(\frac{x \frac{dy}{dx}-y}{y^2}\right) \]

OpenStudy (shamil98):

\[\huge y' = \frac{ -x }{ y }\] \[\huge y''= \frac{ -y + xy' }{ y^2 }\] Plug in y'. \[\huge y'' = \frac{ -y + \frac{ -x^2 }{ y } }{ y^2 }\] \[\huge y'' = \frac{ -y }{ y^2 } + (\frac{ -x^2 }{ y } * y^2)\] \[\huge y'' = \frac{ -1 }{ y } -x^2y\]

OpenStudy (luigi0210):